Two identical balls each have a mass of 10g. What charge should these balls be given so that their interaction equalises the force of universal gravitation acting between them? The radii of the balls may be ignored in comparision to distance between them?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

To exactly cancel the gravitaitonal pull by electrostation repulsion,

\dfrac{Gm^2}{d^2}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q^2}{d^2}
\implies q=m\sqrt{4\pi\epsilon_0 G}=8.6\times 10^{-13}C